OpenStudy (anonymous):

Find the length and width of a rectangle with an area of 2x^2 + x - 3. l = 2x + 3; w = x + 1 l = 2x + 3; w = x - 1 l = 3x + 2; w = x - 1

OpenStudy (anonymous):

@Abdulhameed

OpenStudy (triciaal):

area of a rectangle is length x width option 1 factor the equation the width has the lower value option 2 for each of your answer choices do the multiplication of length x width and the one that gives the area in the question is the correct answer answer b in either case

OpenStudy (anonymous):

(2x+3)(x-1) l = 2x + 3; w = x - 1 the second one

OpenStudy (anonymous):

@xo_kansasprincess_xo sorry for not replying it was 3 a.m here when u mentioned me timezone lol